OpenStudy (idkwut):

True or false: d/dx sqrt(1+e^(3x)) = (3/[2(sqrt1+e^(3x)])^e^(3x)

zepdrix (zepdrix):

\[\Large\bf\sf \frac{d}{dx}\sqrt{1+e^{3x}}\quad=\quad \left(\frac{3}{2\sqrt{1+e^{3x}}}\right)^{e^{3x}}\]What..? Did you make a typo with a ^ in there somewhere maybe?

OpenStudy (idkwut):

That was the option given. I think it's false. There are other options too but I just wanted to check that one so that I can rule it out. Because the answer is actually (3^e^(3x)/(2(sqrt1+e^3x). I think they were trying to trick us with that one but the other options are written in a different form.

zepdrix (zepdrix):

Oh ok :o yah it's false.